August 24th 2022

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • HCAL
      • taking overnight runs
      • adjusting delay on trigger, and latency
      • 40 bins 160 ns
      • delay for trigger adjusted
  • CDet
    • Magnetic shield test
    • finishing efficiency analysis
    • VXS crate
      • need give hostnae cdet crate to Bryan
    • GEM
      • SBS setup - finished - issue taking long runs - MPD firmware had to
      • taking cosmics but issue with scintillators : 3 out of 6 scintillators were working, but issue fixed - had to increase voltage by 50 V - maybe because of HV modules - cables.
      • will take cosmics during night time
      • can try to integrate into full DAQ - need a copy cosmics - need to plug TS to TI
      • pedestal width same as in EEL
      • SBS magnet off : some electrons in GEM, for adjusting timing -
      • BB cabling for trigger and clock : ribbon cable on V1495 was disconnnected
    • hodoscope
      • swapped HV : EPICS looks ok
      • one good cosmic run
      • data looks fine
      • hodoscope ready to go
    • Helicity : toolset for shift crew - need to know how often to start run for feebdack
      • delayed helicity ? Can check if can do in time - last time